### v1

# iRAG: An Incremental Retrieval Augmented Generation System for Videos

Abstract
Retrieval augmented generation (RAG) systems combine the strengths of language generation and information retrieval to power many real-world applications like chatbots. Use of RAG for combined understanding of multimodal data such as text, images and videos is appealing but two critical limitations exist: one-time, upfront capture of all content in large multimodal data as text descriptions entails high processing times, and not all information in the rich multimodal data is typically in the text descriptions. Since the user queries are not known apriori, developing a system for multimodal to text conversion and interactive querying of multimodal data is challenging.
To address these limitations, we propose iRAG, which augments RAG with a novel incremental workflow to enable interactive querying of large corpus of multimodal data. Unlike traditional RAG, iRAG quickly indexes large repositories of multimodal data, and in the incremental workflow, it uses the index to opportunistically extract more details from select portions of the multimodal data to retrieve context relevant to an interactive user query. Such an incremental workflow avoids long multimodal to text conversion times, overcomes information loss issues by doing on-demand query-specific extraction of details in multimodal data, and ensures high quality of responses to interactive user queries that are often not known apriori. To the best of our knowledge, iRAG is the first system to augment RAG with an incremental workflow to support efficient interactive querying of large, real-world multimodal data. Experimental results on real-world long videos demonstrate 23x to 25x faster video to text ingestion, while ensuring that quality of responses to interactive user queries is comparable to responses from a traditional RAG where all video data is converted to text upfront before any querying.
Index Terms:
Video Analytics, Retrieval Augmented Generation (RAG), Large Language Model, Generative AI, Multimodal Data, Natural Language Processing
I Introduction
Many applications routinely collect and store videos for offline analysis. For example, surveillance systems ensure public safety by monitoring public spaces [1], airports, transportation hubs, and critical infrastructure using video cameras. They also store video feeds for offline analysis and investigations. There is also an increasing usage of video analytics for patient monitoring in hospitals and healthcare facilities. Analyzing video feeds from traffic cameras for traffic monitoring, congestion management [2], and incident detection also requires an understanding of information in videos.
Large language models (LLMs) like ChatGPT [3] have demonstrated their proficiency in human-like text-based dialogues. Inspired by the success of LLMs in understanding large repositories of text, recent proposals attempt to understand information in videos by using LLMs [4, 5]. These proposals use vision AI models to analyze content in each video frame, and these models output the visual information in a frame as a text description. For example, an object detector AI model can identify objects in each frame as cars, trucks, bicycles, persons etc., and also output the location of these objects in a frame as human understandable text. Such AI models can process long videos and capture visual information as a text document, which is a concatenation of text descriptions from each frame. LLMs can leverage these text documents using advanced techniques like retrieval augmented generation (RAG) [6], and provide useful responses to interactive queries about the content in the videos.
Prior proposals to understand video content through interactive queries have two major limitations. First, the time required to process entire long videos and generate text descriptions by using complex AI models can be prohibitive. For example, it could take more than a day for complex AI models to analyze visual information in a 24-hour surveillance video and generate text descriptions. This prevents the police from timely analysis of criminal incidents in the video. Second, not all visual information in the videos is captured in the text descriptions, even if we use state of the art complex vision AI models [7, 8, 9, 10]. Furthermore, the user queries are not known apriori, and it is difficult to determine which AI models should be used to convert visual content into text.
To address the critical limitations of prior efforts to interactively query content in long videos, we propose iRAG. Unlike prior proposals [4, 5] that employ many AI models to extract textual information upfront, iRAG uses lightweight AI models to quickly prepare an index for the content in the long videos. During preprocessing of the long video, our primary goal is to extract enough information from the video to be able to index the video data. Such an index facilitates effective extraction and retrieval of more detailed information from select sections of the video in response to interactive user queries.
While utilizing lightweight AI models reduces the indexing time, the text descriptions may not be as rich, and many user queries may not be answerable by an LLM due to lack of facts in the text descriptions. To tackle this challenge, we propose an incremental workflow. When the information extracted during indexing is inadequate to successfully respond to a user query, additional models are used on-demand to extract more details from selective parts of the long video. This strategy ensures that iRAG extracts necessary detailed textual information on-demand, addressing specific queries as they arise, rather than collecting all textual information upfront without knowing the specific details needed to answer arbitrary user queries.
In summary, we make the following contributions:
  1. 1.
    We propose a novel incremental RAG system called iRAG to quickly start interactive querying of large video repositories. Unlike traditional methods that employ many AI models up front to convert entire video information into text, iRAG makes query-aware selection of AI models as necessary. Compared to prior work, iRAG improves the preprocessing time by 23x to 25x, and uses incremental extraction to ensure comparable useful LLM responses.
  2. 2.
    iRAG consists of three novel components that work together to enable incremental RAG: (a) Given a user query, the Query Planner identifies sections of the video that should be analyzed in more detail. The Planner also identifies the AI models that should be used for detailed analysis. (b) The Indexer uses a novel AI-based re-ranking method to significantly reduce the portions of the long video that should be considered for detailed analysis. (c) Finally, the Extractor extracts detailed information, and updates the index to ensure query-relevant retrieval of context that is useful to elicit useful responses from an LLM.
II Motivation
Prior work [4, 5] has augmented user prompts with additional context from stored videos to guide an LLM to produce more contextually relevant and informed responses. These systems produce text descriptions of content in long videos upfront by using a variety of AI models, and they use traditional RAG systems to support interactive querying. We highlight two key limitations that arise when existing methods are applied to real-world videos and scenarios, and we propose a novel, incremental workflow to address them.
II-A Prior work to leverage context in video
Prior work [4, 5] first produces a textual description of information in the entire video. One or more AI models like object detection analyze every frame in the stored video and produce a textual description like a list of objects in the frame. Concatenation of descriptions for each frame results in a long document. Then, a traditional RAG system is used to augment the user prompt with query-specific context in the long document. The augmented prompt guides the LLM to produce more useful responses. We briefly review this process.
Preprocessing: As shown in Figure 1, a long video is divided into smaller clips. Each clip has a pre-determined number of frames. Each frame in a clip is analyzed by one or more vision AI models like object detection to produce a textual description. Text from all the frames in the video is concatenated into a long document. A variety of vision or generative AI models can be used to extract information from each video frame. However, video to text conversion can take a long time depending on the available computational resources, and complexity of the AI models.
Figure 1: Video to text conversion
RAG: The long text document is stored in a vector database. This makes it easy to retrieve portions of the document that are most relevant to a given user query. The long text document is chopped up into chunks where a chunk can be a paragraph, or a sentence or some pre-determined number of consecutive characters. Each chunk is embedded into a vector of pre-determined length by an embedding algorithm, and these vectors are all stored in a vector database. All this makes it easy to identify chunks in the document that are similar to a user query.
For example, given a user query, we can embed the query (by using the embedding model) into a vector. Then, we can compute the dot product of the query vector with each vector in the database, and easily determine the top $k$ vectors in the database that are most similar to the query vector. As shown in Figure 2, $k$ is a configuration parameter of the baseline RAG system.
Since we know the chunk that corresponds to each vector in the database, it is easy to retrieve the $k$ chunks in the long document that are most similar to the user query. The semantic search block in Figure 2 retrieves the relevant query-specific context from the long document. Then, user query is augmented with query-specific context, and the augmented prompt results in a more contextually accurate response from an LLM.
In this paper, we use the preprocessing and retrieval methods shown in Figures 1 and  2 as our baseline to evaluate our proposed new approach.
Figure 2: Conventional RAG
II-B Limitations of Prior work
We identify two critical limitations that make it difficult to use of prior work for advanced understanding of real-world videos.
Long pre-processing time: Since the user queries are not known apriori, all available vision AI models are used to glean information in the stored video. Often, models that can capture more details and nuances in the video are often very computationally expensive. Preprocessing time matters. For example, crimes are time-sensitive, and the sooner law enforcement can analyze surveillance footage, the greater the likelihood of capturing critical evidence while it is still fresh. As we show later, preprocessing of a 24-hour surveillance video using prior work can easily take more than a day. This limits how quickly the context in the videos can be used to solve real-world problems, especially when the timely analysis of surveillance video is crucial for law enforcement (to gather actionable intelligence, identify suspects, or conduct thorough investigations) to solvd crimes and ensure public safety.
Figure 3 shows the preprocessing time bottleneck for querying and analyzing long videos from several traffic intersections in New York city [11]. We consider three videos that are 1, 7, or 30 days long. We use the DETR object detection model [7] and the GRiT dense captioning model [10] to extract text information from the videos. The average per-frame latency times are approximately 70 ms, and 1500 ms for DETR and GRiT, respectively, on a server equipped with an AMD Ryzen 5950X 16-Core processor and a NVIDIA GeForce RTX 3090 GPU. Although we process only one frame in every one-second video clip, the video-to-text generation process for for 1-day, 7-day, and 30-day videos requires approximately 1.57 days, 11 days, and 46.4 days, respectively. Such long video-to-text conversion times hinder development of practical systems to quickly query and understand long videos from real-world applications.
1 day7 days30 days0020202020404040401.571.571.571.571111111146.446.446.446.4Preprocessing Time (days)
Figure 3: Time to generate a long document from real-world videos
Information loss: During the conversion of video (vision domain) to text (text domain), there is a loss of information due to the shift in modalities. There are many details in the video that are not captured in the long text document, even if we use a lot of AI models during preprocessing. Since the user queries are not known apriori, often the long text document lacks sufficient information that can help the LLM provide a useful response to a user query. When this happens, there is no mechanism in prior approaches to go back to the stored videos and efficiently extract query-specific additional information (perhaps, by using new models that were not used or available during preprocessing).
Our new approach: Instead of executing all vision AI models upfront to extract textual information from entire videos, we propose an incremental RAG (iRAG) system that does query-aware, on-demand extraction of text data from select portions of the videos. Furthermore, iRAG automatically selects appropriate vision AI models that are necessary during the incremental workflow to answer the user query.
Such a capability is not present in prior methods [4, 5] that solely rely on long preprocessing times to convert video to text once before any interactive querying of the long videos.
III Incremental RAG
III-A Preprocessing in iRAG
Unlike prior approaches, the goal of preprocessing in iRAG is to quickly build an index for the long video. For example, a lightweight object detection model like DETR [7] can analyze each frame in the video and identify objects, their types, their location and size (bounding box). Such information forms a natural index to quickly identify the clips in the video that may have information relevant to a user query. Frames provide a natural timeline, with each frame containing information that can be searched to identify relevant clips for responding to a user query.
For example, to answer the query “Is there a FedEx truck in a long video?”, the index (i.e. the object detection information) can be used to quickly identify the relevant clips where a truck was detected. Subsequent detailed extraction of text information from these select clips (by using an AI model that can pick up text in images) can confirm if and where a FedEX truck is seen in the long video. We remark that a variety of indexes can be quickly constructed by using lightweight vision AI models. Also, the indexes do not have to be human-readable or understandable, and in the future, it is likely that the features used to construct the index are automatically generated by AI models. However, for ease of understanding of our new approach, we limit the discussion to the use of two popular lightweight models (DETR [7] and CLIP [12]) during preprocessing in iRAG. Such an index is human understandable, and it provides transparency and explainability of the query responses from iRAG.
III-B Query-Response
Given a user query, iRAG retrieves the relevant context from the information collected during preprocessing by using the conventional RAG approach. Then, we construct a suitable prompt for the LLM by using the following prompt template:
Prompt 1
You operate as a chatbot that is supported by a retrieval augmented generation system. You will utilize the given context and your knowledge to answer queries. If you are unable to answer a query, your response is “Unable to answer query. Please run additional models”.
Context: {context}
Query: {query}
We invoke the incremental flow when the LLM is unable to answer a user query. Next, we describe the iRAG system, and its key components.
III-C System Overview
Figure 4 shows an overview of the proposed iRAG system. The incremental flow (which is invoked only when the LLM is unable to answer a query) is shown using dotted lines.
There are three key components that support the query-aware, on-demand incremental flow: query planner, indexer and extractor.
Figure 4: System overview of iRAG. The dotted lines represent the additional workflow beyond a conventional RAG.
Our Query Planner retrieves query-aware relevant context from the information gathered in preprocessing, and recommends AI models that should be used to extract more details from selected clips in the long video. These additional details are necessary to satisfactorily answer the user query. Context retrieved by the Planner consists of two parts: the top-$N$ chunks in TextDB (containing information from the DETR model) that are most similar to the user query, and the top-$F$ frames in ImageDB (containing information from the CLIP model) that are most similar to the user query. Note that $N$ and $F$ are configuration parameters for the iRAG system.
Our Indexer further refines (and reduces) the context from the Planner to produce at most $k$ video clips that will be analyzed the Extractor. These video clips are the context clips. As shown in Figure 4, $k$ is also a configuration parameter for the iRAG system. Unlike the the interpretation of $k$ in the baseline system (where $k$ chunks of text context are used to construct the prompt for the LLM), iRAG interprets $k$ to be the upper limit on the number of context clips that will be sent to the Extractor. Since detailed extraction of information from the context clips dominates the query processing time, the choice of $k$ affects the quality of interactive querying experience. Furthermore, the configuration parameters $N$ and $F$ are typically an order or magnitude larger than $k$ (as described in the Experimental Results section IV).
Our Extractor extracts more details from the context clips by using AI models recommended by the Planner. The text and image vectors obtained from the extraction of context clips are added to the TextDB and ImageDB, respectively, to update the index. We then re-run the user query by retrieving updated context, and the likelihood of eliciting a useful response from the LLM increases.
The iRAG system can be configured to run the incremental flow in a loop for a preset maximum number of times, after which the system reports that the user query cannot be answered. In practice, a single iteration of the incremental flow is adequate to successfully answer most queries (as observed in the results reported in the Experimental Results section IV).
The response to a query consists of two parts: a textual response to the user query, and a set of video clips from the long video that support the textual response (refer to Figure 11 for an illustration). This allows the user to detect incorrect responses (or hallucinations) from the LLM. In the future, we can also construct automatic response verifiers that determine whether the human understandable textual response and video clips support each other.
III-D Query Planner
Figure 5 shows the workflow in our planner. Given a query, the Planner computes a query embedding vector by using the same embedding algorithm that the RAG system used to embed preprocessing text chunks into vectors in TextDB, and retrieves the top-$N$ vectors in TextDB that are most similar to the query embedding vector. Chunks that correspond to the top-$N$ vectors are the top-$N$ chunks, which are the most relevant chunks for the user query.
Figure 5: Query planner workflow
Improved context retrieval: To improve the quality of the retrieved context, we use an additional vector database ImageDB to store frame vectors. Lightweight AI models like CLIP embed visual information in each frame of the long video into a CLIP embedding vector, which is also the frame vector. All the frame vectors produced during preprocessing are stored in ImageDB. The Planner computes the CLIP embedding vector for the user query, and retrieves the top-$F$ frame vectors that are most similar to the user query vector. The frames corresponding to the top-$F$ vectors are the top-$F$ frames.
Context from the Planner includes two parts: Top-$N$ context chunks from TextDB, and the top-$F$ context frames from the ImageDB. As we show next, the Indexer does further processing to significantly reduce this context to at most $k$ context clips.
Figure 6: Indexer workflow
III-E Indexer
The Indexer refines the context from the Planner, and ensures that we have at most $k$ context clips for the Extractor.
As stated earlier, $k$ is a configuration parameter of the iRAG system. A large $k$ implies a higher limit on the context clips, and the detailed extraction of so many clips will take time and adversely affect the quality of interactive querying. On the other hand, a very small $k$ can result in too few clips for detailed extraction, and the updated index may not have the context necessary for the LLM to provide a useful response to the user query. So, choice of $k$ is critical to ensure a balance between interactivity and usefulness of LLM responses.
Merge: As shown in Figure 6, the top-$N$ chunks from the Planner are further refined using a re-ranking algorithm (described next) to produce fewer context chunks. The merge component in the Indexer merges the top-$F$ frames and the clips corresponding to the context chunks by concatenation to produce at most $k$ context clips for the Extractor.
If the re-ranking algorithm is not used, then the clips corresponding to the top-$N$ chunks are merged with the top-$F$ frames to create at most $k$ context clips. However, if our new re-ranking algorithm is used, then a significant reduction in the top-$N$ clips is possible.
Re-ranking: A KNN classifier (described next) is used to filter the top-$N$ chunks into two classes: “accept” or “reject”. Figure 7 shows the re-ranking workflow. Each vector in the top-$N$ vectors is concatenated with the query vector, and the KNN classifier assigns an “accept” or “reject” label to the concatenated vector. Any top-$N$ vector that has the “reject” label is eliminated from further consideration. The remaining vectors form a new ordered list. The chunks corresponding to these vectors also form an ordered list, which is provided as input to the merge block.
Training of KNN classifier: We train a classifier by using data from our baseline system, and iRAG’s Planner. Given a query, and a long video, we analyze the context retrieved by the baseline system, and iRAG. If a chunk retrieved by iRAG matches a chunk retrieved by the baseline system, then we assign a label of “1” (accept) to the iRAG chunk. Otherwise, we assign a label of “0” (reject) to the iRAG chunk. This creates labeled data that associates the concatenation of a user query vector and an iRAG chunk vector with a label “1” or label “0”. By using the test queries in VQA-2 dataset, we created labeled training data that was used to train our KNN classifier offline. After training, our KNN classifier assigns a label “1” to an unlabeled iRAG context chunk if 3 out of 5 neighbors of the concatenated query-chunk vector also have the label “1”.
Figure 7: Re-ranking workflow
III-F Extractor
The Extractor receives the context clips from the Indexer. It also receives (from the Planner) a list of AI models to run on each clip. The text data obtained after detailed extraction of the context clips is used to update TextDB. Then, the user query is processed again to create a prompt for the LLM. Figure 8 shows how the Planner, Indexer and Extractor work together in the iRAG system to incrementally update TextDB.
Figure 8: Interaction of key components in iRAG
IV Experimental results
We compare iRAG with the baseline video RAG system described in Section III. Baseline uses GRiT [10] for preprocessing, while iRAG uses DETR [7] and CLIP [12]. iRAG employs GRiT when preprocessed information cannot answer a user query. We compare the pre-processing time, average query processing time, and accuracy of the two systems. We briefly describe our video datasets in Table I, and the queries we pose on these datasets. We use outputs from the baseline system as the ground truth.
IV-A Datasets
VQA-v2 validation dataset [13]: This dataset has 40,504 images. Of the 214,354 queries that are posed on these images, we selected 7799 queries, all of which start with “Is there”. We concatenate the images corresponding to these queries to create a long video (about 2 hours and 10 minutes long at 1 frame per second). We also randomly choose 1000 queries from the 7799 queries. These 1000 queries are used to evaluate iRAG and the baseline system.
MSR-VTT [14]: This dataset has small clips of YouTube videos, with a caption per video. We concatenate several video clips to create a long video (about three and half hours long at 20fps). This dataset also has 318 queries.
StreetAware: [11]: This dataset has 7.75 hours of video (2 TB of raw data) of traffic activity at four New York City traffic intersections. We selected a 46-minute video for our evaluation. This dataset does not have queries. As described later, we generated appropriate queries by suitably prompting an LLM.
Tokyo MODI [15]: This dataset is a video of a busy street in Tokyo. Like others, we also used 2 hours of this video [15]. This dataset has no queries, and we prompt an LLM to generate queries for our evaluation.
TABLE I: Datasets used for evaluating iRAG
Video Dataset Type Duration Keyframes # of test queries
VQA-v2 Traditional 02:09:59 7799 1000
MSRVTT Traditional 03:29:35 6291 318
StreetAware Real-world 00:46:45 2363 40
Tokyo MODI Real-world 02:00:23 1444 14
IV-B Queries
Although two of our datasets (VQA-v2 and MSRVTT) come with queries, the real world datasets like StreetAware [11] and MODI [15] do not. Therefore, we use an LLM to generate appropriate queries. For this purpose, we ran the GRiT model [10] on both videos to generate captions, which are then used as context for the following prompt:
Prompt 2
Given a context, generate a question for the context. Here are a few examples of questions:
- Is there a white van?
- Is there a person with a backpack?
- Is there a bus?
Return the question without any additional text. Think you are an investigator querying a textual description of a video.
Context: {context}
By repeatedly prompting the LLM with various contexts, we generated 40 queries for the StreetAware dataset and 16 queries for the Tokyo MODI datasets.
IV-C System implementation details
We implemented iRAG using the Langchain framework [16]. The vector database (with the FAISS index [17]) is from Langchain, and the LLM is OpenAI’s gpt-3.5-turbo model [18] API. We used DETR [7], an object detector, to do a fast preprocessing of videos for iRAG. In the incremental flow, we utilized GRiT [10], a popular image caption generator. However, this model requires significantly more computational resources (averaging approximately 1500 ms per frame) compared to the DETR object detector (averaging approximately 70 ms per frame). We conducted experiments on a server equipped with an AMD Ryzen 5950X 16-Core processor and an NVIDIA GeForce RTX 3090 GPU.
IV-D Evaluation metric
For the baseline and iRAG, we retrieve $k$ chunks as context ($k$ is a pre-determined configuration parameter for RAG). These chunks correspond to video clips. Given a query, we compute the ratio of the number of video clips common to both systems and the number of video clips in the baseline. If the ratio is 1, then both iRAG and the baseline retrieve the same video clips as context for the query. We define recall@k as the average of these ratios for all queries in a dataset. By varying $k$, we can compute different recall@k and study the impact of $k$ on the quality of context retrieved by iRAG (compared to the baseline).
Table II shows the context quality recall@k for different values of $k$. For example, for the VQA dataset, when iRAG uses only DETR object detector for indexing (i.e. preprocessing), the recall@k for $k=20$ is $0.815$. This means that across all the queries in the dataset, there was on average about 81% overlap of video clips retrieved as context by iRAG and the baseline. Higher the value for recall@k, the closer are the responses from the LLM for the baseline and iRAG. However, naively increasing $k$ to improve context quality will also lead to an increased processing time in the incremental flow. Increasing $k$ can increase the number of video clips that will require detailed extraction during the incremental flow. This additional processing time can adversely affect interactive querying experience.
IV-E System evaluation
We systematically evaluate the benefits from different enhancements in iRAG, compared to the baseline. We also study the impact of varying $k$ on the context quality and interactivity. This aids in the selection of a suitable $k$ for interactive querying.
Benefit from using imageDB: By using lightweight vision-language models like CLIP, we can extract richer indexes in the preprocessing phase of iRAG. As shown in Table II, the context video clips retrieved by iRAG when both CLIP and DETR models are used in preprocessing have a better context quality than the case where only DETR is used.
Performance of iRAG on the StreetAware and MODI datasets are shown in Figure 10 and Table VII, respectively. For these datasets, we show the human-understandable response video clips, and LLM responses from the iRAG system.
TABLE II: Performance of iRAG on VQA-v2 and MSRVTT datasets.
Dataset System Indexer r@1 r@2 r@4 r@8 r@20
VQA-v2 iRAG DETR 0.2564 0.3593 0.4961 0.6368 0.8153
DETR + CLIP 0.3299 0.4127 0.5314 0.6607 0.8251
MSRVTT iRAG DETR 0.4098 0.5135 0.6652 0.8003 0.9223
DETR + CLIP 0.4127 0.5206 0.6716 0.8028 0.9240
Benefit from chunk re-ranking: Results in Table II do not consider the case where the chunks retrieved by the Planner module are re-ranked by the Indexer to eliminate the less important chunks. The benefit from re-ranking context chunks is shown in Table III and Table IV. For the VQA-v2 dataset, with $k=8$, the Planner module retrieves 8 context chunks and the recall score is 66.07%. However, with chunk re-ranking, based on the query, the Indexer is able to eliminate less important chunks from the proposed 50 chunks from the Query Planner. Thus, the average number of chunks selected by the Indexer is only 6.88 while the recall score is 74.14%. Across all the 1000 queries for this data set, the average number of clips sent for detailed extraction in the incremental flow was only 6.88 chunks rather than 8 chunks. This results in reduced processing times for detailed extraction with accuracy improvement for the iRAG system with the refinement module. Therefore, iRAG suggests a cost-effective approach to extracting data for a query. For the classifier training, we select 1000 query from the training data of VQA-v2, for each query we run the GRiT model to extarct ground truth clips. Then we label the each concatenated query-chunk embedding vector as 1 if the chunk contains clips from ground truth, other wise give a 0. Then we train KNN classifier on this. After the KNN is trained, we test it on 1000 queries from the validation samples from VQA-v2 dataset to report results.
TABLE III: Benefit from re-ranking (VQA-v2 dataset)
Method initial k avg. # of chunks per query recall score
iRAG [w/o re-ranking] 8 8 0.6607
iRAG [w re-ranking] 50 6.88 0.7414
As the increase of k indicates the increase in total query processing time due to extracting unnecessary details using the Extractor, we define this k as a cost metric. We observe that for a specific k value (cost), the iRAG with re-ranking module offers improved accuracy (recall@k) score compared to iRAG w/o the re-ranking module. To visualize this, for iRAG, in Table IV, we set the initial k that the query planner offers are 10-50 chunk proposals. If we follow the query planner proposals, we may result in extracting frames from 10-50 chunks by the Extractor. Our iRAG with re-ranking module refines the k value and select it on average from $1.49\sim 6.88$ chunks. Now, if we observe the recall@k score for similar value of k ranging from 2-8 in iRAG w/o re-ranking, we observe that the iRAG with re-ranking improves the recall score by $1.9\%\sim 13.8\%$. Hence, the iRAG with re-ranking module is cost-effective and efficient.
TABLE IV: Impact of re-ranking on context quality
iRAG iRAG
w/o re-ranking with re-ranking
cost recall@k cost (k) initial k recall@k
(k) [after re-ranking] [before re-ranking]
2 0.367 1.49 10 0.386
2.84 20 0.528
4 0.501 4.17 30 0.639
6 0.567 5.49 40 0.683
8 0.624 6.88 50 0.741
Preprocessing time:
Preprocessing of the long videos using DETR is a lot faster than preprocessing with GRiT. Table V shows that iRAG’s preprocessing time is a 23x to 25x less than the preprocessing time of the baseline.
TABLE V: Preprocessing time of iRAG and baseline
Data System Model Preprocessing Time
VQA-v2 iRAG DETR + CLIP 48 min 29 sec
RAG GRiT 1093 min 36 sec
MSRVTT iRAG DETR + CLIP 8 min 23 sec
RAG GRiT 199 min 50 sec
Query processing time vs $k$: Query processing time is important. It determines if the query-response session is interactive. It includes retrieval of the context clips by the Planner, optimizations in the Indexer, and detailed extraction of video clips by the Extractor. Figure 9 shows how interactivity (the query processing time) varies with $k$.
There are two important characteristics of iRAG. First, from the inset that shows the query processing time for the first 200 queries, we observe that the query processing time for the first few queries is quite high. However, the query processing time drops dramatically as we process more queries. This is because the information from the detailed extraction of clips for the first few queries is added to the text and image vector DBs, and subsequent queries do not trigger a lot of detailed extraction. Second, as we increase $k$, the number of context clips increase and the incremental extraction time increases, leading to an increase in query processing time. For example, the query processing time for $k=2$ for the first few queries is approximately 25-30 seconds. However, for the same queries, the query processing time jumps to 70-75 seconds when $k=4$. However, irrespective of the value of $k$, as shown in the inset for queries 200 - 1000, query processing time drops rapidly to less than 5 seconds as we process more queries. We also re-ordered the 1000 queries a few times, and observed that the above observations hold.
Figure 9: Query processing time distribution for different values of $k$
Re-ranking context clips in the indexer eliminates less important clips, and improves interactivity. This is because due to re-ranking, not all the $k$ context clips are sent for detailed extraction. Re-ranking effectively reduces $k$ for the Extractor module, and the reduced detailed extraction processing time leads to significantly improved interactivity.
Note that the LLM query execution time remains almost constant across queries for both iRAG and the baseline. As a result, we have excluded this time from the query processing time calculations.
TABLE VI: Detailed extraction statistics
k
Fraction of
video extracted
(VQA-v2)
Fraction of
video extracted
2 0.437 0.323
4 0.578 0.392
6 0.670 0.428
8 0.724 0.445
10 0.771 0.458
20 0.907 0.493
30 0.968 0.503
40 0.986 0.508
50 0.995 0.509
Incremental extraction:
The baseline does detailed extraction of all the video clips during preprocessing. In contrast, iRAG does opportunistic, query-driven detailed extraction of clips as necessary. Table VI shows the fraction of the clips for which the Extractor performed detailed extraction for different values of $k$. For example, after processing all the queries in MSRVTT dataset, and even for large values of $k$, the Extractor performed detailed extraction for only about half the video clips.
For VQA dataset, iRAG processes about 77% of the images for reasonable values of $k$ (between 2 and 10). The percentage is a higher because the long video was artifically created by us. Here, several images are packed into a video clip, and we do detailed extraction on clips rather than images. If we treat a video clip to be a frame, then the fraction of images processed by the Extractor is similar to the MSRVTT dataset case.
IV-F Human Evaluation
Figure 10 shows frames in the context clips retrieved by iRAG and the baseline for a variety of queries. Visual inspection also shows that the context retrieved in iRAG is similar to the context retrieved by the baseline, except for one example. For the query, “Is there a white van?”, iRAG context is a poor match with the context from the baseline.
Table VII shows the LLM responses to queries in the Tokyo MODI dataset. In almost all cases, the responses from iRAG align well with the responses from the baseline system, except for the query “Are there any neon signs in Chinese in the video?”. The iRAG system does find clips with neon signs though. However, more measures are required to detect LLM hallucinations [19] and re-issue the user query.
Figure 10: Qualitative evaluation of iRAG system over StreetAware dataset.
TABLE VII: LLM API response on MODI street video dataset using iRAG
Query Incremental RAG (iRAG) RAG
Were there any black cars on the road? Based on the given context, there was a black car on the road at the time 1:52:10. Based on the given context, there was a black car on the road at 1:36:05. Therefore, the answer to the question “Were there any black cars on the road?” is yes, there was a black car on the road at 1:36:05.
Were there any people riding bicycles? Based on the given context, there is evidence of a person riding a bicycle in the following segment: ”When 1:37:15 I found the word shop on the building, people walking on the street, people crossing the street, the photo was taken in black and white, the people are walking, person walking across the street, white van parked on the side of the road, a large billboard, person wearing a jacket, person walking on sidewalk, a green building, person crossing the street, two people walking across the street, person on a bicycle, car parked on the street, person crossing the street.”Therefore, the answer to the question ”Were there any people riding bicycles?” is ”Yes, there was a person riding a bicycle.” Yes, there were people riding bicycles. In the context provided, it mentions ”person riding a bike” at both 1:58:15 and 1:12:50. Therefore, the answer to the question ”Were there any people riding bicycles?” is yes.
Were there any green trees lining the sidewalk? Based on the given context, I can confirm that there were green trees lining the sidewalk in the following instances: - When 1:52:10- When 0:59:30 - When 1:11:15 - When 1:44:30 - When 2:$0$ When 0:42:20 - When 0:29:45- When 0:10:35. Therefore, the answer to the question “Were there any green trees lining the sidewalk?” is “Yes, there were green trees lining the sidewalk in multiple instances.” Yes, there were green trees lining the sidewalk in both instances. The context provided mentions “green trees lining sidewalk” in both the timestamps 1:55:10 and 1:35:20. Therefore, the answer to the question ”Were there any green trees lining the sidewalk?” is yes.
Were there any neon signs in Chinese visible in the video? I can not answer. Yes, there were neon signs in Chinese visible in the video. The context provided mentions ”I found a neon sign in chinese” at both 0:28:50 and 0:33:35. Therefore, the answer to the question is ”Yes, there were neon signs in Chinese visible in the video.”
What color is the car on the street? Based on the given context, the color of the car on the street is black. This information is found in the context for the time stamp 1:52:10, where it states ”I found black car on the street.” Therefore, the color of the car on the street is black. The color of the car on the street is green and white. This information is found in the context provided, specifically in the segment ”When 1:28:25”. Therefore, the answer to the question ”What color is the car on the street?” is green and white.
Were there any people wearing hats? Based on the given context, I will search for instances of people wearing hats in the provided images. Let me analyze the images to find the answer to your question. Yes, there were people wearing hats in the given context. In the first timestamp (1:57:55), it mentions ”man wearing a black hat” and in the second timestamp (1:19:50), it mentions ”woman wearing a black hat.” Therefore, the answer to the question ”Were there any people wearing hats?” is yes.
Were there any green plants in the video? Based on the given context, the answer to the question ”Were there any green plants in the video?” is ”Yes, there were green plants in the video.” This is evident from the following excerpts from the context:- ”a green building” - ”a green tree” - ”a bunch of green plants” - ”a green tree in a parking lot” - ”green trees lining sidewalk”- ”a bunch of green produce” - ”a bunch of green flowers” - ”a planter with green leaves” - ”a green tree with lots of leaves” - ”a green tree in a planter” - ”a plant with purple flowers” - ”a planter with a lot of flowers” These excerpts provide clear evidence that there were green plants in the video. Yes, there were green plants in the video. At 1:03:55, green plants were found in a planter and green trees were lining the sidewalk. At 1:01:35, green plants were found in white vases.
Were there any green trees with lots of leaves? Based on the given context, there are multiple instances of green trees with lots of leaves mentioned. Therefore, the answer to the question ”Were there any green trees with lots of leaves?” is ”Yes.” Yes, there were green trees with lots of leaves mentioned in the given context. The context specifically states ”I found a green tree with lots of leaves” in both instances at 0:19:35 and 0:15:35. Therefore, the answer to the question ”Were there any green trees with lots of leaves?” is yes.
Were there any umbrellas seen on the street? Based on the provided context, there is no mention of umbrellas being seen on the street. Therefore, the response to the question ”Were there any umbrellas seen on the street?” is ”No, there were no umbrellas seen on the street.” Based on the given context, there were umbrellas seen on the street. At 1:01:50, the umbrellas are described as green, and at 1:28:15, an umbrella is described as pink. Therefore, the answer to the question ”Were there any umbrellas seen on the street?” is yes, there were umbrellas seen on the street.
Were there any bicycles in the video? Based on the given context, I can not answer. Step 1: Analyze the given context for any mention of bicycles. Step 2: Review the context for any descriptions or sightings of a person riding a bike.Step 3: Identify any specific details related to bicycles, such as the presence of a bike or someone riding a bike. Step 4: Based on the analysis of the context, provide a response to the question. Response: Yes, there were bicycles in the video. At 0:05:40, a person was found riding a bike, and at 1:12:50, another person was found riding a bike.
Were there any green buses in the video? I can not answer. Yes, there were green buses in the video. At 0:35:30, the context mentions “green bus on the street” and “white writing on the bus.” Additionally, at 0:37:05, it states “the bus is green.” Therefore, based on the given context, we can confirm that there were green buses in the video.
Were there any people wearing checkered shirts? I can not answer. Based on the given context, there were no mentions of people wearing checkered shirts. Therefore, the response to the question ”Were there any people wearing checkered shirts?” is “I can not answer”.
Were there any green traffic lights in the video? I can not answer. Based on the given context, there is no mention of any green traffic lights in the video. Therefore, the answer to the question “Were there any green traffic lights in the video?” is “No, there were no green traffic lights in the video.”
Were there any people wearing white jackets? Based on the given context, there is evidence of a person wearing a white jacket in the following segment: “When 1:54:30. I found red white and blue bus, a sign on a building, white van on street, person wearing a black hat, woman wearing a white jacket, black car on the street, the word shop on the building, the people in the city, a box of umbrellas, people walking on the street, person wearing a black jacket, green trees lining sidewalk.” Therefore, the answer to the question ”Were there any people wearing white jackets?” is “Yes.” Yes, there were people wearing white jackets. In the given context, it is mentioned that “woman wearing white jacket” was found at 1:45:35 and “person wearing white jacket” was found at 0:43:05. Therefore, the answer to the question “Were there any people wearing white jackets?” is yes.
V Prototype Implementation
Figure 11: Prototype implementation of the video iRAG system.
We have developed a prototype that allows users to interact with the iRAG system. Figure 11 illustrates the user interface for the iRAG system. Initially, the user uploads a video and starts the preprocessing stage by clicking on “Start Preprocessing,” utilizing fast AI models. Once preprocessing is complete, the iRAG system is ready to process the query.
In this example, the DETR object detection model is employed for preprocessing. Now, for detection-based queries such as “Is there a bus?”, the indexer efficiently indexes the video and provides the query’s answer. For more advanced user queries, like “What is the color of the bus?”, the indexer suggests the necessary video clips. The Extractor then uses the GRiT model to extract detailed text information from these clips. Subsequently, the extracted text serves as context, and the LLM API is invoked along with the query to generate the final response.
VI Related Work
Traditional Video Understanding Models. Traditional video understanding models have focused on designing foundational models for specific tasks [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. These models distinguished themselves by incorporating a variety of neural network architectures and employing diverse training methodologies. These models demonstrate superior performance when analyzing short video clips lasting several seconds, but they encounter difficulty when tasked with processing longer videos.
Video-Language Models. Researchers have made significant progress in the field of multimodal representation learning [30, 31, 32], particularly emphasizing the convergence of video and language understanding [33, 34, 35]. They employ a two-stage approach, commencing with the pretraining of a foundational video-language model using extensive pairs of video and text, and subsequently fine-tuning on specific downstream tasks. Similarly, recent studies propose the integration of a frozen Language-Modeling Module (LLM) alongside an image encoder and a limited set of learnable modules, introducing a technique termed “visual instruction tuning” [36, 37]. The main goal of this technique is to fine-tune the learnable modules, aiming to enhance the LLMs’ capacity to generate textual descriptions for video content. Although all these models exhibit promising performance, they may face challenges when dealing with long-duration videos. Notably, these models are typically trained on short video clips, potentially introducing difficulties in understanding longer video sequences.
Video question answering system [38, 39, 40] is extensively studied in the literature. Unfortunately, most of them mainly focus on short video clips of images. In recent works, the focus has been on understanding complex relationships in long videos, such as cause-and-effect and sequential order [41, 42]. Typically, questions in long-form video question-answering systems require reasoning about multiple events, various levels of detail, and causality. For each query, with efficient segmentation of videos, the whole work might be repetitive for searching the corresponding relevant frames which is almost impossible for long videos.
LLMs in Text-Based Video Understanding. Large language models (LLMs) like ChatGPT have exhibited notable proficiency in handling text-based dialogues, and effectively managing input and output in textual format. Recently, LLMs connect visual foundation models that create a new dimension for advanced video understanding [43, 44]. Inspired by the success of reasoning and interaction with LLMs, recent studies have redefined the challenge of understanding long video content through a question-and-answer framework [4, 5, 35]. Initially, these methods segment a lengthy video into smaller clips and utilize readily available vision and vision-language models to convert visual data from each clip into text. Subsequently, they employ LLMs to generate textual responses to queries associated with the video, utilizing textual information as context. In contrast to these methods, which typically rely on a set of models to extract textual information from videos upfront, our approach aims to execute models incrementally based on queries, extracting the necessary textual information on-demand to address each specific query.
VII Conclusion
We introduced a novel incremental Retrieval-Augmented Generation (iRAG) system for enhancing understanding of multimodal data like videos and images.
Instead of using many AI models to upfront extract text data from the entire corpus of multimodal data, iRAG first quickly indexes the corpus of multimodal data by using lightweight AI models. There will be user queries for which the information extracted from the multimodal data during index building is insufficient. In such cases,
iRAG invokes the incremental workflow to extract more detailed information from select, query relevant multimodal data. Such new information is used to retrieve an updated context, which results in a useful response to the user query. We evaluated iRAG for interactive video understanding, but the iRAG system is useful for combined understanding of a variety of other multimodal data like audio, lidar and time-series. The AI models required to capture content in such multimodal data as text, or quickly build indexes, are different. However, like video, such multimodal data also has time as a marker, and the multimodal content is easily embedded into vectors, making iRAG indispensable for advanced, interactive understanding of multimodal data.